In 3 spins of an 8 spot spinner, what is the probability of getting exactly one number that is 5 or less than 7.
Answer by stanbon(75887)

In 3 spins of an 8 spot spinner, what is the probability of getting exactly one number that is 5 or less than 7.
-----------------
P(get 5 or 1-6) = P(get 5) + P(get 1-6) - P(5 AND 1-6)
-----
Note: P(5 AND 1-6) = P(get 5)
----------------------------
So, P(get 5 or 1-6) = P(get 1-6) = 6/8 = 3/4
================================================
Comment: I may not be reading your problem correctly
but that's what it says to me.
---------------------------------------
Cheers,
Stan H.